# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v1

@@ V Discussion
Unpaired pre-training of the content/style model boosts the practical applicability of PnP-MUNIT, e.g. in the case of LUMC-TRA where the T1W/T2W data imbalance is considerable. In this work, we limited our baselines to those algorithms that closely resemble ours in terms of design or data requirement. As future work, a more comprehensive study could be conducted with a broader range of methods and evaluation criteria, e.g. comparing generalizability across accelerations and sampling patterns with guided unrolled networks, given a fixed budget of paired training data. At reconstruction time, we assumed the reference and reconstruction images to be spatially aligned. The CR update, which can correct minor registration errors implicitly, would likely break down at the typically observed levels of patient motion. A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further relaxing the overall data constraints.
In terms of reconstruction quality, we observed (Section IV-B) that using both measured k-space and reference scan via PnP-MUNIT was more beneficial than (a) using either one of them and (b) combining both in a naive way or based on hand-crafted prior, suggesting that our approach maximally exploits the complementary information. Compared to PnP-CNN, PnP-MUNIT allowed up to 32.6% more acceleration for a given SSIM. [^1] Based on linear interpolation of LUMC-TRA SSIM, PnP-MUNIT and PnP-CNN allowed $R$=10 and $R$=6.7, respectively, the maximum difference in $R$, at median SSIM=0.906. Hence, the maximum time saving 32.6%. In the radiological task (Section IV-C), PnP-MUNIT produced diagnostic-quality reconstructions at $R$=3, a difference of 33.3% over the clinical reconstructions of $R$=2.222We define diagnostic quality as visibility of at least 90% of the metastases and scoring at least “fairly diagnostic quality” on all other criteria. In clinical practice, while 10% of the metastases may not be detected in the reconstructed T2W scan, they would be prominently visible in the corresponding gadolinium-enhanced T1W scan, and hence would not go undetected. However, it was somewhat worse than PnP-CNN, while theoretically, the additional guidance should result in at least equivalent image quality. We argue that the likely cause is the limitations of the dataset and the specific content/style model, not of the proposed general approach. First, in LUMC-TRA, T1W scans had 2-3 times lower resolution than T2W scans, limiting the value of guidance (Section IV-A1). In the long run, this may be solved by co-designing a higher-resolution reference sequence with the target sequence to optimize guidance. Second, and more importantly, correcting for content discrepancy using CR relies on computing gradients through the decoder $G_{2}$. If $G_{2}$ is highly non-linear and the reference content is not sufficiently close to the optimum, CR may converge to a sub-optimal point leading to texture artifacts in the reconstruction, as seen in Fig. 6 at $R$=4. In the radiological study, precisely this phenomenon is likely to have impacted the visual and diagnostic quality of PnP-MUNIT. Hence, to improve clinical utility, the content/style model must be improved using specialized architecture and regularization.
VI Conclusion
@@
In this work, we introduced PnP-MUNIT, a novel approach to multi-contrast reconstruction combining content/style modeling with iterative reconstruction, which is interpretable in that it allows the two stages to be analyzed separately. By exploring the interpretability aspect, two limiting factors of guided reconstruction were identified – the amount of effective content in a two-contrast dataset and content discrepancy. On real-world clinical data, PnP-MUNIT provided up to 32.6% more acceleration over PnP-CNN for a given SSIM, enabling sharper reconstructions at high accelerations. Given the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-MUNIT produced diagnostic-quality images at $R$=3, enabling 33.3% more acceleration over clinical reconstructions. To progress towards a practical implementation, future work will focus on improving the content/style model and reconstruction latency and incorporating online registration into the reconstruction.

### v2

@@ VI Discussion
In terms of training data requirements, PnP-MUNIT can make use of the larger amounts of the image data available and can be applied to situations where end-to-end methods are infeasible since it does not rely on k-space training data. Moreover, unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-MUNIT, e.g. in the case of LUMC-TRA where the T1W/T2W data imbalance was considerable. As indicated in our NYU DICOM benchmark, PnP-MUNIT performs similarly or better than state-of-the-art end-to-end methods. In our LUMC multi-coil benchmark, we limited our baselines to those algorithms that were feasible given the data constraint. As future work, a more comprehensive study should be conducted with a broader range of methods and evaluation criteria, e.g. comparing generalizability across accelerations and sampling patterns with multi-contrast unrolled networks, given a fixed budget of paired training data. Contrary to the training stage, we strictly assumed the spatial alignment of the reference and target images in the reconstruction stage. The CR update, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion. A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further improving practical applicability.
In terms of reconstruction quality, we observed in Section V-C that the true added value of PnP-MUNIT was at the high acceleration factors ($R$=8 and $R$=10) where the k-space data is scarce and the reference information becomes more valuable. However, the risk of model hallucinations increases at these accelerations, raising concerns about the accuracy of the anatomy represented in the image. Given the advantages of content/style decomposition, it may be possible to leverage the contrast-independence property of the content to automatically detect local hallucinations in the reconstruction and to define a corrective process to minimize it. This is another topic for future work. At lower accelerations, on the other hand, we observed that PnP-MUNIT often produced slightly lower metrics compared to PnP-CNN, e.g. in LUMC-COR. This was also observed in the radiological evaluation on LUMC-TRA samples. This is counter-intuitive at first glance since PnP-MUNIT has access to additional side-information and should, in principle, perform at least as good as PnP-CNN. However, this can be explained by the fact that at lower accelerations, the content/style model becomes the limiting factor for reconstruction quality. This is because unlike the CNN denoiser (Eq. (4)) of PnP-CNN, the content consistency operator (Eq. (14)) of PnP-MUNIT radically changes the image and hence, this operation is sensitive to the model’s overall performance. The model performance is in turn influenced by the training dataset and the network architecture. First, regarding data, the T1W scans in LUMC datasets had an in-plane resolution 1.3-2.5 times lower than that of the T2W scans, limiting the quality of the model (as indicated by its low optimal content capacity) and the overall value of guidance. In the long run, this may be solved by using 3D reference and target sequences of matched-resolution to optimize guidance. Second, regarding model design, correcting for content discrepancy using the CR module relies on computing gradients through the decoder $G_{2}$ (Eq. (16)). If $G_{2}$ is highly non-linear and the reference content is not sufficiently close to the optimum, CR may converge to a sub-optimal point leading to texture artifacts in the reconstruction, as seen in Fig. 10 at $R$=4. In our proof-of-concept, we used the same general model architecture proposed in [22]. Better designs that are more conducive to CR may exist. Hence, future efforts must considerably focus on improving the content/style model using specialized architecture and regularization.
VII Conclusion
@@
In this work, we introduced PnP-MUNIT, a modular approach to multi-contrast reconstruction combining content/style modeling with iterative reconstruction, which offers the reconstruction quality of end-to-end methods under stronger training data constraints. At its core is the content consistency operation, which provides regularization at the level of the image’s semantic content. We defined two quantities that determine the efficacy of this operation, namely optimal content capacity and content discrepancy, and provided several ways of maximizing this efficacy. On real-world clinical data, PnP-MUNIT provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations. In the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-MUNIT produced diagnostic-quality images at $R$=3, enabling 33.3% more acceleration over clinical reconstructions. To progress towards practical implementation of our proof-of-concept, future work will focus mainly on improving the content/style model and reconstruction latency, tackling the problem of model hallucinations, and incorporating online registration into the reconstruction.